OpenStudy (anonymous):

What is the equation of the line that passes through the point (2, 3) and is perpendicular to the line y = -x - 3? Choose one answer. a. y = -x - 1 b. y = -x + 5 c. y = x + 1 d. y = x + 5

OpenStudy (asnaseer):

first, do you know what the slope of the line y = -x - 3 is?

OpenStudy (anonymous):

No is it -1 not really sure

OpenStudy (asnaseer):

yes its -1 next, you should know that the slopes of perpendicular lines multiply to give -1. so if the line y = -x - 3 has slope -1, then what do you think will be the slope of a line perpendicular to this?

OpenStudy (anonymous):

would it be 1

OpenStudy (asnaseer):

correct. finally, the slope-intercept form of a line is written as: y = mx + c where m is the slope and c is the y-intercept. now you know the slope is 1, so you can write the equation of the perpendicular line as: y = x + c all that remains is to put in the coordinate (2, 3) to calculate c. so put in x=2, y=3 into this equation to work out c and you have then solved the question.

OpenStudy (anonymous):

I don't understand sorry I did 3=2+c is that right so far

OpenStudy (asnaseer):

that is correct, now find the value for c that satisfies this equation.

OpenStudy (anonymous):

so I got c=1

OpenStudy (asnaseer):

perfect, so the answer to your question would be?

OpenStudy (anonymous):

y=-x-1?

OpenStudy (asnaseer):

we ended up with: y = x + c and you found c=1, so answer is?

OpenStudy (anonymous):

y=x+1?

OpenStudy (asnaseer):

thats it!

OpenStudy (anonymous):

Thanks so much!

OpenStudy (asnaseer):

yw